# Effects of domain walls and chiral supercurrent in quantum anomalous Hall Josephson junctions

### v2

@@ III Numerical results
III.1 Current-phase relation
To start with, we investigate the current-phase relation with varying the chemical potential $\mu$ in the pure JJ, as shown in Figs. 2(a) and (b). The Josephson current is computed employing the recursive Green-function method as proposed in Refs. M1 ; M2 ; M3 . In Fig. 2(a), the current-phase relation of $I_{s}/I_{c}$ exhibits a period of $2\pi$ and displays a typical sinusoidal shape when $\mu=0,0.1,0.21$ with a magnetic flux $\phi=0$. Note that it slightly deviates from the sinusoidal relation when $\mu=0.4,0.7,1$. Notably, upon introducing $\Phi$ to the QAHI region, a magnetic field-controlled $0-\pi$ transition is observed at $\mu=0.21$ with $\phi=0.0024$ in Fig. 2(b).
Figure 2: The normalized supercurrent $I_{s}/I_{c}$, where $I_{c}$ is the critical Josephson current, changes with (a) the superconducting phase difference $\varphi$ when $\phi=0$, and (b) when $\phi=0.0024$, where $\phi$ denotes the magnetic flux on each lattice. Different colored curves correspond to different values of the chemical potential $\mu$. (c) The QAHI energy spectrum with a length of $L=40$ is calculated by using a lattice model described by Eq. 2. $\mu_{1}=0$ and $\mu_{2}=0.21$ are the chemical potentials, indicated by the green dashed lines. Density of states of (d) $\mu_{1}=0$ and (e) $\mu_{2}=0.21$ in the QAHI region. (f) The supercurrent $I_{s}$ with the superconducting phase difference $\varphi=\pi/2$ as a function of the total magnetic flux of the central region $\Phi$ for the chemical potential $\mu=0.21$. In all cases, we take $M_{z}=0.3$ and the temperature $T=\Delta/200$.
Figure 3: Quantum interference patterns in different regimes. The critical Josephson current $I_{c}$ versus the magnetic flux $\Phi$ in the unit of $\Phi_{0}$. (a) CARs happen when only chiral edge states are included with $\mu=0$. $I_{c}$ exhibits the unconventional AB oscillation characterized by a periodicity of $2\Phi_{0}$. The inset showcases the dependence of $I_{c}$ on the aspect ratio $L/W$ (depicted as red circles), and a fitting curve is provided as a solid black line. We take $M_{z}=0.3$. (b)-(e) The behavior of $I_{c}$ deviates from the $2\Phi_{0}$-periodic oscillation as $\mu$ takes on values of 0.1, 0.21, 0.4, and 0.7, respectively. (c) The wavefunctions at the left and right edges exhibit overlapping behavior, making the occurrence of LARs at a single edge. $I_{c}$ is notably marked by a $\Phi_{0}$ periodicity. It is noteworthy that when $\mu>0.21$, bulk carriers actively engage in the transport process. (d) The emergence of central and side lobes is a consequence of the interference pattern transitioning towards a regime dominated by bulk carriers. (f) $I_{c}$ displays the typical asymmetric Fraunhofer pattern with $\mu=1$. In this regime, the bulk carriers uniformly flow through the whole QAHI region. In all cases, the blue (orange) solid lines represent $I_{c}$ curves corresponding to $M_{z}=\pm 0.3$.
To grasp the origin of this transition, we examine the density of states in the QAHI region corresponding to the Fermi energies $\mu_{1}=0$ and $\mu_{2}=0.21$, depicted in Figs. 2(d) and (e). When $\mu_{1}=0$, the states localize at the left and right QAHI-SC interfaces without any presence of bulk carriers in the QAHI region, resulting solely in CARs. Conversely, by shifting $\mu$ to $0.21$, the density of states distinctly shows that the states at both the left and right QAHI-SC interfaces broaden and spatially overlap. This broadening allows the supercurrent to flow across the QAHI bulk region, resulting in the occurrence of LARs occur along with the CARs. In this case, this supercurrent under the influence of magnetic flux exhibits behavior akin to the Fraunhofer pattern, expressed as $I_{s}\propto\frac{\rm{sin}(\pi\Phi/\Phi_{0})}{\pi\Phi/\Phi_{0}}\rm{sin}(\varphi)$. Consequently, at certain magnetic flux values, a $0-\pi$ transition may occur. To support this statement, we plot the supercurrent $I_{s}$ as a function of $\Phi$ for $\varphi=\pi/2$, fitting the data points to the ideal Fraunhofer pattern, illustrated in Fig. 2(f). The observed slight deviation from the ideal Fraunhofer pattern fit may stem from the inhomogeneous carrier density within the QAHI region. Therefore, we conclude that the $0-\pi$ transition underscores the crossover from the chiral-edge to bulk dominant transport.
Figure 4: The spatial distribution of local critical current in the QAHI region [(a), (c), and (e)] and at the QAHI-SC interface [(b),(d), and (f)] is examined under varying chemical potentials: $\mu=$0 [(a),(b)], 0.21 [(c),(d)], and 1 [(e),(f)]. In all cases, $M_{z}=0.3$, $\Phi=0$, and the temperature $T=\Delta/200$.
III.2 Three distinct quantum interference patterns
Next, we examine the quantum interference patterns within the QAHI-based junction. At a chemical potential of $\mu=0$, as depicted in Fig. 3(a), the QAHI region exclusively accommodates chiral edge carriers. At the interface between the QAHI and the SC, an electron from one chiral edge mode can tunnel into a hole on the opposite edge through the CAR process. Similarly, the holes are then reflected as electrons along the other QAHI-SC interface. This entire process forms a loop, completing the transfer of the supercurrent, and leading to an intriguing AB oscillation pattern. This interference pattern, a hallmark of chiral edge states, exhibits two distinctive features: (1) Unlike the conventional $\Phi_{0}$ periodicity, the critical Josephson current $I_{c}$ has a periodicity of $2\Phi_{0}$. (2) The minima of $I_{c}$ are non-zero. These findings align with prior research studies f1 ; f2 ; f3 ; f4 ; f5 ; f6 ; f7 . We note that the exponential decay of the peak value of $I_{c}$ with the aspect ratio $L/W$ is suitably described by the equation $I_{c}\propto\rm{exp}(-\alpha L/W)$ (see the inset of Fig. 3(a)), where $\alpha$ is a fitting constant. Furthermore, by shifting the chemical potential to $\mu=0.1$, an observed phase shift arises from the finite chemical potential. This occurrence is tied to the sustenance of the Josephson current by Andreev bound states forming a closed loop. Consequently, the product of the Fermi wavelength and the circumference, $k_{F}*L_{d}$, introduces a phase shift, resulting in current oscillations relative to the chemical potential. Interestingly, adjusting $\mu$ to $0.21$ leads to a substantial overlap of chiral edge mode wavefunctions at both the left and right QAHI-SC interfaces due to their considerable broadening. This configuration triggers LARs at a single edge and within the inner QAHI region, giving rise to a $\Phi_{0}$ periodicity in $I_{c}$ (see Fig. 1(c)) KTLaw .
Figure 5: (a) Schematic diagram of a QAHI-based JJ with a random DW structure (top view). The DW size is denoted by $d_{1}\times d_{2}$. The quantum interference patterns are affected by the different strengths of magnetic disorder $W$ when (b) $\mu=0$ and (c) $\mu=1$. In all cases, the DW size is $d_{1}\times d_{2}=10\times 4$, and the DW probability $P=0.5$. We average over 50 random DW configurations.
@@ III.2 Three distinct quantum interference patterns
To trace the evolution of the Josephson current from edge states to the bulk, we increase the chemical potential to $\mu>0.21$, disrupting the previously observed $2\Phi_{0}$ or $\Phi_{0}$ periodicity. Interestingly, the Fraunhofer pattern with a distinct central peak of width $2\Phi_{0}$ forms when $\mu>0.4$ (Figs. 3(d)-(f)), signaling the increased dominance of bulk states in the transport process. This occurrence correlates with the population of bulk carriers within the QAHI region, leading to a consistent appearance of LARs at QAHI-SC interfaces as $\mu$ increases (refer to Fig. 1(d)). Due to the increase in bulk carriers, a noticeable enhancement in the maxima of $I_{c}$ is also observed, as seen in Figs. 3(a)-(f), exceeding hundreds of times. Furthermore, the Fraunhofer pattern displays an asymmetric feature due to broken time-reversal symmetry, namely, $I_{c}(M_{z},\Phi)\neq I_{c}(M_{z},-\Phi)$ when the magnetic field direction is reversed. Besides, the inversion symmetry invariance of $I_{c}$ leads to the relationship $I_{c}(M_{z},\Phi)=I_{c}(-M_{z},-\Phi)$ S1 ; S2 . In the subsequent calculations, we focus exclusively on the scenario where $M_{z}=0.3$ due to the aforementioned relationship. In summary, $I_{c}$ of the QAHI-based JJs can exhibit three distinct quantum interference patterns by tuning the chemical potential $\mu$: the unconventional $2\Phi_{0}$-periodic oscillation, the $\Phi_{0}$-periodic oscillation and the asymmetric Fraunhofer pattern.
Figure 6: The spatial distribution of local critical current in the QAHI region [(a) and (c)] and at the QAHI-SC interface [(b) and (d)] affected by the DW structure is examined under varying chemical potential values: $\mu$=0 [(a),(b)] and 1 [(c),(d)]. In this case, the DW size is $d_{1}\times d_{2}=10\times 4$, the magnetic disorder strength $W=0.8$, and the DW probability $P=0.5$. In all cases, $M_{z}=0.3$, $\Phi=0$, and the temperature $T=\Delta/200$.
To further reveal the origin of different interference patterns, we investigate the spatial distribution of local critical current $j_{c}$ in the QAHI region M1 ; M2 ; M3 . The local critical current is usually used to represent the current distribution of various systems, e.g., quantum Hall systems LHL . Figure 4 demonstrates the local current along the x-direction, $j_{c}(y)$, corresponding to the superconducting phase difference $\varphi$, precisely when the total current $\sum\limits_{y=1}^{W}j_{c}(y)$ reaches its maximum. At $\mu=0$, the current predominantly flows along the QAHI region’s two edges, as evidenced in the spatial distribution of $j_{c}$ displayed in Figs. 4(a) and (b). This observation aligns with the phenomenon depicted in Fig. 1(b), where transport within the central region is solely governed by the propagation of chiral edge carriers. Additionally, this current distribution is consistent with the quantum interference pattern shown in Fig. 3(a), characterized by a periodicity of $2\Phi_{0}$. Upon increasing $\mu$ to 0.21, the edge current widens noticeably at the left and right QAHI-SC interfaces as shown in Figs. 4(c) and (d). This visibly demonstrates the overlapping wavefunctions at these interfaces. Consequently, local Andreev pairs emerge both on a single edge and within the inner QAHI region (see Fig. 1(c)), consistent with the scenario depicted in Fig. 3(c). Moreover, Figs. 4(e) and (f) show a bulk-dominated transport process, as the $\mu$ value is further increased to 1. Thus, local Andreev pairs become uniformly distributed across the QAHI region (see Fig. 1(d)), resulting in a Fraunhofer pattern, as observed in Fig. 3(f).
III.3 Effects of domain walls
@@
Figure 7: The quantum interference patterns are affected by the DW structure at $\mu=0$ in a larger magnetic field range with different block sizes $d_{1}\times d_{2}$: (a) $50\times 20$, (b) $20\times 10$, and (c) $10\times 4$. In all cases, the magnetic disorder strength $W=1$ and the DW probability $P=0.5$. We average over 15,25,50 random DW configurations in (a),(b), and (c), respectively.
@@
Given the inevitable appearance of DWs in magnetically doped topological insulators, we utilize a percolation model to simulate the effects of random DWs on quantum interference patterns Li . Fig. 5(a) demonstrates the division of the central region into multiple blocks, each consisting of specific unit cells sized $d_{1}\times d_{2}$. The magnetization within each block is randomly chosen to potentially flip with a probability of $P$, while the flipped magnetization strength is uniformly distributed in the range of $[-W/2,W/2]$ with the disorder strength $W$. The fragility of the $\Phi_{0}$-periodic oscillation pattern to the influence of DWs leads us to exclude its detailed discussion in this section.
@@
In Figs. 5(b) and (c), we display the DW-averaged critical Josephson current, denoted as $\overline{I}_{c}$. Initially, focusing on the case with $\mu=0$ in Fig. 5(b), it is observed that $\overline{I}_{c}$ remains relatively stable (in comparison with Fig. 3(a)) when $W=0.2$, indicating the robustness of chiral edge states against weak magnetic disorder ($W$). Then, at $W=0.5$, the partial destruction of chiral edge modes diminishes the peak of the critical Josephson current $\overline{I}_{c}$. Despite DWs causing the emergence of some bulk carriers in the QAHI region, the $2\Phi_{0}$ periodicity remains notably stable, owing to the dominance of transport by chiral edge modes. After that, increasing $W$ to 0.8 initiates two simultaneous processes: the breakdown of chiral edge modes and a comparable contribution of bulk carriers to that of chiral edge carriers. Consequently, both the $2\Phi_{0}$ periodicity and the characteristic shape of the quantum interference pattern are lost. Conversely, at $\mu=1$, we observe remarkable stability in the interference pattern of bulk carriers, maintaining the shape of the asymmetric Fraunhofer pattern even at $W=0.8$. This suggests that the unconventional $2\Phi_{0}$-periodic oscillation is more susceptible to disruption compared to the resilient nature of the asymmetric Fraunhofer pattern under DW structures. Furthermore, the values of $\overline{I}_{c}$ gradually decrease with increasing $W$ in Figs. 5(b) and (c), manifesting the influence of disorder effects.
Figure 8: The averaged critical Josephson currents $\overline{I}_{c}$ vary with the DW probability $P$ with $\mu=0$, $W=0.5$, and $d_{1}\times d_{2}=20\times 10$. In all cases, we average over 25 random DW configurations.
@@
Next, we examine the spatial distribution of the local critical current $j_{c}$ within the QAHI region and at the QAHI-SC interface in Fig. 6, considering the existence of DW structures. In our calculation, we maintain a constant total magnetic flux $\Phi=0$ and compute the local current $j_{c}(y)$ along the x-direction when the total current $\sum\limits_{y=1}^{W}j_{c}(y)$ reaches its maximum. Due to the presence of DW structures, the local critical current $j_{c}$ exhibits static eddy-like currents in addition to constant currents. In Figs. 6(a) and (b) at $\mu=0$, it is evident that the chiral edge modes are disrupted, and some bulk carriers emerge within the inner QAHI region due to the random distribution of DWs across the QAHI region. This observation aligns with the behavior depicted in the curve corresponding to $W=0.8$ in Fig. 5(b). Conversely, when $\mu=1$, the distribution of $j_{c}$ exhibits characteristics akin to typical bulk transport, facilitating the stability of the asymmetric Fraunhofer pattern within DW structures. Notably, in Figs. 6(c) and (d), $j_{c}$ values oscillate more prominently compared to Figs. 4(c) and (d).
@@
To gain deeper insights into how DWs affect the edge current, we explore the quantum interference pattern at $\mu=0$ by varying the sizes of DWs under a strong magnetic disorder strength $W=1$, spanning a wider range of magnetic fields. In Fig. 7(a), an emergent central lobe is observed within the interval $\Phi/\Phi_{0}\in(-2,2)$, alongside side lobes at higher values of $\Phi$. Notably, the average critical current $\overline{I}_{c}$ displays significant fluctuations, which can be attributed to the presence of DWs. Additionally, the increase in the maxima of the central lobe is linked to the emergence of bulk carriers caused by a limited number of DWs, particularly when the DW size is $d_{1}\times d_{2}=50\times 20$. This interference pattern remarkably resembles to the Fraunhofer pattern illustrated in Fig. 3(f). The emergence of bulk carriers in the QAHI region due to the presence of DWs is the underlying cause of this phenomenon. When we decrease the DW size to $d_{1}\times d_{2}=20\times 10$ to allow for a larger quantity of DWs, a more distinct Fraunhofer-like pattern is evident within the $\overline{I}_{c}$ values in Fig. 7(b). Again, the presence of bulk carriers contributes to the enhancement of maxima of the central lobe, compared to the scenario depicted in Fig. 3(a). When the DW size is further decreased to $d_{1}\times d_{2}=10\times 4$, as depicted in Fig. 7(c), the increased visibility of the periodicity in the higher lobes reveals a distinct Fraunhofer-like pattern. In particular, the periods of the central and side lobes are twice those of the conventional Fraunhofer pattern. To conclude, the observed Fraunhofer-like pattern in chiral JJs likely extends beyond our model and is prevalent in those doped with magnetic impurities, holding promise for further experimental verification.
At last, we briefly discuss another factor affecting the behavior of the quantum interference pattern in DW structures: the DW probability $P$, as illustrated in Fig. 8. With an increase in $P$, more DWs emerge in the QAHI region, leading to an evolution of the interference pattern toward a Fraunhofer-like pattern, aligning with our previous analysis. In summary, the interference pattern within the DW structure can be influenced by factors such as magnetic disorder strength $W$, DW size $d_{1}\times d_{2}$, and DW probability $P$.
IV conclusion
@@
In this work, we explore the transport properties of the chiral JJs based on the QAHIs connecting to two s-wave superconducting electrodes, through numerical calculation of the Josephson current ($I_{c}$). We present a systematic crossover from edge-state to bulk-state dominated supercurrents, with a notable $0-\pi$ transition that becomes apparent with adjustments in the chemical potentials under non-zero magnetic flux. This transition shows the competition between bulk and chiral edge transport. Moreover, we find an evolution among three distinct quantum interference patterns through adjustments in the chemical potentials: from a $2\Phi_{0}$-periodic oscillation pattern, to a $\Phi_{0}$-periodic oscillation pattern, and then to an asymmetric Fraunhofer pattern. Finally, a Fraunhofer-like pattern, with periods twice those of the conventional one, emerges due to the coexistence of the chiral edge state and the domain wall states, even when the chemical potential resides within the gap. This phenomenon is likely widespread in chiral JJs doped with magnetic impurities. Our findings also hold promise for direct experimental verification.

### v3

@@ III Numerical results
III.1 Evolution of quantum interference patterns
[…]
@@ III.1 Evolution of quantum interference patterns
To trace the evolution of the Josephson current from edge states to the bulk further, we continue to increase the chemical potential to $\mu>0.42\Delta$, disrupting the previously observed $2\Phi_{0}$ periodicity. Interestingly, the Fraunhofer pattern with a distinct central peak of width $2\Phi_{0}$ forms when $\mu>0.8\Delta$ (Figs. 2(d)-(f)), signaling the increased dominance of bulk states in the transport process. This occurrence correlates with the population of bulk carriers within the QAHI region, leading to a consistent appearance of LARs at QAHI-SC interfaces as $\mu$ increases (refer to Fig. 1(d)). Due to the increase in bulk carriers, a noticeable enhancement in the maxima of $I_{c}$ is also observed, as seen in Figs. 2(a)-(f), exceeding hundreds of times. Furthermore, the Fraunhofer pattern displays an asymmetric feature due to broken time-reversal symmetry, namely, $I_{c}(M_{z},\Phi)\neq I_{c}(M_{z},-\Phi)$ when the magnetic field direction is reversed. Besides, the inversion symmetry invariance of the QAHI’s Hamiltonian $H(\mathbf{k})$ leads to the relationship $I_{c}(M_{z},\Phi)=I_{c}(-M_{z},-\Phi)$ S1 ; S2 . The inversion symmetry holds because $\mathcal{P}H(-\mathbf{k})\mathcal{P}^{-1}=H(\mathbf{k})$ with $\mathcal{P}=\sigma_{z}\tau_{0}$, where $\sigma_{z}$ and $\tau_{0}$ are the Pauli matrix and the unit matrix in spin space and layer space, respectively. In the subsequent calculations, we focus exclusively on the scenario where $M_{z}=0.3$ due to the aforementioned relationship between $I_{c}(M_{z})$ and $I_{c}(-M_{z})$. In summary, $I_{c}$ of the QAHI-based JJs undergoes a transition from a $2\Phi_{0}$-periodic oscillation pattern to an asymmetric Fraunhofer pattern by tuning the chemical potential $\mu$. Note that we do not discuss the case where the chemical potentials $\mu$ and $\mu_{s}$ are different here because the periodicity of $2\Phi_{0}$ remains consistent when the difference is not large (See Appendix A for more information).
@@
Figure 4: (a) Schematic diagram of a QAHI-based JJ with a random magnetic domain structure (top view). The domain size is denoted by $d_{1}\times d_{2}$. The quantum interference patterns are affected by the different strengths of magnetic disorder $w$ when (b) $\mu=0$ and (c) $\mu=2\Delta$. In all cases, the domain size is $d_{1}\times d_{2}=10a\times 4a$, and the domain probability $P=0.5$. The remaining parameters are set as: $M_{z}=0.6\Delta$, $\mu_{s}=\mu$, and the temperature $T=\Delta/200$. We average over 50 random domain configurations Ave .
To further reveal the origin of different interference patterns, we investigate the distribution of current $j_{c}$ in QAHI region when the total current has its maximum value M1 ; M2 ; M3 . The local critical current is usually used to represent the current distribution of various systems, e.g., quantum Hall systems LHL . Figure 3 demonstrates the distribution of current along the x-direction, $j_{c}(y)$, corresponding to the superconducting phase difference $\varphi$, precisely when the total current $\sum\limits_{y=1}^{W}j_{c}(y)$ reaches its maximum. At $\mu=0$, the current predominantly flows along the QAHI region’s two edges, as evidenced in the spatial distribution of $j_{c}$ displayed in Figs. 3(a) and (b). This observation aligns with the phenomenon depicted in Fig. 1(c), where transport within the central region is solely governed by the propagation of chiral edge carriers. Additionally, this distribution of current is consistent with the quantum interference pattern shown in Fig. 2(a), characterized by a periodicity of $2\Phi_{0}$. Moreover, Figs. 3(c) and (d) show a bulk-dominated transport process, as the $\mu$ value is increased to $2\Delta$. Thus, local Andreev pairs become uniformly distributed across the QAHI region (see Fig. 1(d)), resulting in a Fraunhofer pattern, as observed in Fig. 2(f).
III.2 Effects of magnetic domains
Given the inevitable appearance of domains in magnetically doped topological insulators, we utilize a percolation model to simulate the effects of random domains on quantum interference patterns Li . In realistic samples, domains typically exhibit irregular shapes. However, the specific shapes of domains are not of particular importance to our study. The bulk carriers induced by domains that are essential to the physics discussed in this section. Therefore, for the sake of simplicity, we model domains as rectangular in shape.
Figure 5: The spatial distribution of current in the QAHI region [(a) and (c)] and at the QAHI-SC interface [(b) and (d)] when the total current has its maximum value affected by the domain structure is examined under varying chemical potential values: $\mu$=0 [(a),(b)] and $2\Delta$ [(c),(d)]. In this case, the domain size is $d_{1}\times d_{2}=10a\times 4a$, the magnetic disorder strength $w=1.6\Delta$, and the domain probability $P=0.5$. In all cases, $M_{z}=0.6\Delta$, $\Phi=0$, $\mu_{s}=\mu$, and the temperature $T=\Delta/200$.
@@ III.2 Effects of magnetic domains
Fig. 4(a) demonstrates the division of the central region into multiple blocks, each consisting of specific unit cells sized $d_{1}\times d_{2}$. The magnetization within each block is randomly chosen to potentially flip with a probability of $P$, while the flipped magnetization strength is uniformly distributed in the range of $[-w/2,w/2]$ with the disorder strength $w$.
@@
In Figs. 4(b) and (c), we display the domain-averaged critical Josephson current, denoted as $\overline{I}_{c}$. Initially, focusing on the case with $\mu=0$ in Fig. 4(b), it is observed that $\overline{I}_{c}$ remains relatively stable (in comparison with Fig. 2(a)) when $w=0.4\Delta$, indicating the robustness of chiral edge states against weak magnetic disorder ($w$). Then, at $w=\Delta$, the partial destruction of chiral edge modes diminishes the peak of the critical Josephson current $\overline{I}_{c}$. Despite domains causing the emergence of some bulk carriers in the QAHI region, the $2\Phi_{0}$ periodicity remains notably stable, owing to the dominance of transport by chiral edge modes. After that, increasing $w$ to $1.6\Delta$ initiates two simultaneous processes: the breakdown of chiral edge modes and a comparable contribution of bulk carriers to that of chiral edge carriers. Consequently, both the $2\Phi_{0}$ periodicity and the characteristic shape of the quantum interference pattern are lost. Conversely, at $\mu=2\Delta$, we observe remarkable stability in the interference pattern of bulk carriers, maintaining the shape of the asymmetric Fraunhofer pattern even at $w=1.6\Delta$. This suggests that the unconventional $2\Phi_{0}$-periodic oscillation is more susceptible to disruption compared to the resilient nature of the asymmetric Fraunhofer pattern under domain structures. Furthermore, the values of $\overline{I}_{c}$ gradually decrease with increasing $w$ in Figs. 4(b) and (c), manifesting the influence of disorder effects.
Figure 6: The Fraunhofer-like patterns. The quantum interference patterns are affected by the domain structure at $\mu=0$ in a larger magnetic field range with different block sizes $d_{1}\times d_{2}$: (a) $50a\times 20a$, (b) $20a\times 10a$, and (c) $10a\times 4a$. In all cases, the magnetic disorder strength $w=2\Delta$ and the domain probability $P=0.5$. The remaining parameters are set as: $M_{z}=0.6\Delta$, $\mu_{s}=\mu$, and the temperature $T=\Delta/200$. In (a)-(c), the orange regions show the central lobes of the Fraunhofer-like pattern. We average over 15,25,50 random domain configurations in (a),(b), and (c), respectively.
@@
Next, we examine the spatial distribution of current $j_{c}$ within the QAHI region and at the QAHI-SC interface when the total current has its maximum value in Fig. 5, considering the existence of domain structures. In our calculation, we maintain a constant total magnetic flux $\Phi=0$ and compute the distribution of current $j_{c}(y)$ along the x-direction when the total current $\sum\limits_{y=1}^{W}j_{c}(y)$ reaches its maximum. Due to the presence of domain structures, the distribution of current $j_{c}$ exhibits static eddy-like currents in addition to constant currents. In Figs. 5(a) and (b) at $\mu=0$, it is evident that the chiral edge modes are disrupted, and some bulk carriers emerge within the inner QAHI region due to the random distribution of domains across the QAHI region. This observation aligns with the behavior depicted in the curve corresponding to $w=1.6\Delta$ in Fig. 4(b). Conversely, when $\mu=2\Delta$, the distribution of $j_{c}$ exhibits characteristics akin to typical bulk transport, facilitating the stability of the asymmetric Fraunhofer pattern within domain structures. Notably, in Figs. 5(c) and (d), $j_{c}$ values oscillate more prominently compared to Figs. 3(c) and (d).
@@
To gain deeper insights into how domains affect the edge current, we explore the quantum interference pattern at $\mu=0$ by varying the sizes of domains under a strong magnetic disorder strength $w=2\Delta$, spanning a wider range of magnetic fields. In Fig. 6(a), an emergent central lobe is observed within the interval $\Phi/\Phi_{0}\in(-2,2)$, alongside side lobes at higher values of $\Phi$. Additionally, the increase in the maxima of the central lobe is linked to the emergence of bulk carriers caused by a limited number of domains, particularly when the domain size is $d_{1}\times d_{2}=50a\times 20a$. This interference pattern remarkably resembles to the Fraunhofer pattern illustrated in Fig. 2(f). The emergence of bulk carriers in the QAHI region due to the presence of domains is the underlying cause of this phenomenon. When we decrease the domain size to $d_{1}\times d_{2}=20a\times 10a$ to allow for a larger quantity of domains, a more distinct Fraunhofer-like pattern is evident within the $\overline{I}_{c}$ values in Fig. 6(b). Again, the presence of bulk carriers contributes to the enhancement of maxima of the central lobe, compared to the scenario depicted in Fig. 2(a). When the domain size is further decreased to $d_{1}\times d_{2}=10a\times 4a$, as depicted in Fig. 6(c), the increased visibility of the periodicity in the higher lobes reveals a distinct Fraunhofer-like pattern. In particular, the periods of the central and side lobes are twice those of the conventional Fraunhofer pattern. To conclude, the observed Fraunhofer-like pattern in chiral JJs likely extends beyond our model and is prevalent in those doped with magnetic impurities, holding promise for further experimental verification. In summary, the interference pattern within the domain structure can be influenced by factors such as magnetic disorder strength $w$ and domain size $d_{1}\times d_{2}$.
IV discussion and conclusion
While QAHIs can manifest in various material classes, our work focuses exclusively on magnetically doped topological insulators. The lack of comprehensive experimental results on QAHI-based JJs is primarily due to the significant experimental challenges. This challenge arises from the difficulty of achieving coexistence among the QAHI and superconductivity. Specifically, realizing the QAHIs requires magnetic doping within topological insulators. Although magnetic dopants are crucial for realizing QAHI, they also pose challenges for maintaining superconductivity. Therefore, achieving an appropriate magnetic doping ratio is crucial in experimental setups. Additionally, the presence of magnetic dopants raises questions regarding the formation of domains, which is one of the primary aspects addressed in our work. This results in the appear of the anomalous Fraunhofer-like pattern, which is likely widespread in chiral JJs doped with magnetic impurities.
@@
In this work, we explore the transport properties of the chiral JJs based on the QAHIs connecting to two superconducting electrodes, through numerical calculation of the Josephson current ($I_{c}$). We present a systematic transition from edge-state to bulk-state dominated supercurrent as the chemical potential varies. This transition leads to an evolution from a $2\Phi_{0}$-periodic oscillation pattern to an asymmetric Fraunhofer pattern. Furthermore, we observe the emergence of a novel Fraunhofer-like pattern, characterized by periods twice those of the conventional pattern. This emergence is attributed to the coexistence of chiral edge states and bulk states induced by magneitc domains, even when the chemical potential is within the gap. This phenomenon is likely widespread in chiral JJs doped with magnetic impurities. Our findings also hold promise for direct experimental verification.